a)Let f(x)=tan x
i)Show that there exists no c in the interval ]0, pi[such that f'(c)=0 even though f(0)=f(pi)=0
ii)Explain why the result in part (a) does not violate Rolle's Theorem.
1
Expert's answer
2020-05-07T20:20:10-0400

a) "f'(x) = sec^2 x \neq 0" at any "c \in (0,\pi)", even through "f(0)=f(\pi)=0."


b) Rolle's Theorem has three hypotheses:

f is continuous on the closed interval [a,b].

f is differentiable on the open interval (a,b).

f(a)=f(b).

Then there exist at least one "c \in (a,b)" such that "f'(c)=0."


Now given function f(x) = tan(x) is not continuous on "[0,\pi ]," because "f(x)= tan(x)" is not defined at "x = \pi/2."

And The derivative function sec2(x) is not defined at "\pi/2", hence the given finction is not diferentiable in "(0,\pi)."

And "f(0)=f(\pi)=0."

Thus, results are not contradictory to Rolle's theorem. Rolle's theorem violates because in this problem Rolle's theorem hypotheses do not hold.